Changing cash conversion cycle Camp Manufacturing turns over its inventory five times each year, has an average payment period of 35 days, and has an average collection period of 60 days. the firm has annual sales of $3.5 million and cost of goods sold of $2.4 million.

a. Caluclate the firm's operating cycle and cash conversion cycle.

b. What is the dollar value of inventory held by the firm?

c. If the firm could reduce the average age of its inventory from 73 days to 63 days by how much would it reduce its dollar investment in working capital?
Answer #1

daysof inventory=365/inventory turn over=365/5=73

days of sales outstanding = 365/receivable turnover=60

OC= daysof inventory + days of sales outstanding=73+60=133

CCC= days of sales outstanding + days of inv.-days of payment=60+73-35=98

b) avg inv=cogs/inv. turnover=2.4/5= 0.48 million

c)73 days

inv. turnover=365/73=5

63 days

inv. turnover=365/63=5.79

reduction in WC= 5.79-5=0.79 million